OpenStudy (anonymous):

simplify t^2+6t/t-1 - 7/t-1

OpenStudy (anonymous):

(t²+6t)/(t-1) - 7/(t-1)

OpenStudy (lalaly):

(t^2+6t-7)/t-1 (t+7)(t-1)/t-1 = t+7

OpenStudy (anonymous):

(t²+6t)/(t-1) - 7/(t-1) = (t²+6t-7)/(t-1) = ((t-1)(t+7))/(t-1) = (t-7)

OpenStudy (anonymous):

is it that answer or yours?

OpenStudy (lalaly):

t+7

OpenStudy (lalaly):

u got it write... only the final answer is wrong...check it again